Use the sample from Tooic 2 DQ 1 and test that the population mean for rolling a singke die is 3.5.
Null hupothesis- population mean equals 3.5
Alternative hypothesis- population mean can not equal 3.5.

Results for Rolling die 50 times:
1,3,2,6,3,4,4,4,2,3,1,1,5,1,4,6,5,5,1,2,4,3,1,4,5,5,5,2,5,3,1,3,2,4,2,5,5,2,4,4,6,5,3,2,4,2,4,6,1,2
Mean: 3.34
Mode: 4
Median: 3.5
Standard Deviation: 1.634
Variance: 2.632
Answer #1

Mean: 3.34

Standard Deviation: 1.634

.........

Ho :   µ =   3.5
Ha :   µ ╪   3.5
Level of Significance ,    α =    0.05                  
sample std dev ,    s =    1.6340                  
Sample Size ,   n =    50                  
Sample Mean,    x̅ =   3.3400                  
                          
degree of freedom=   DF=n-1=   49                  
                          
Standard Error , SE = s/√n =   1.6340   / √    50   =   0.2311      
t-test statistic= (x̅ - µ )/SE = (   3.340   -   3.5   ) /    0.2311   =   -0.69
                          
  
p-Value   =   0.4920   [Excel formula =t.dist(t-stat,df) ]  

          
Decision:   p-value>α, Do not reject null hypothesis                   

it is concluded that population mean is equal to 3.5

...

thanks

revert back for any doubt